# Simulator Ensembles for Trustworthy Autonomous Driving Testing

### v1

@@ Abstract
Our case study, which involves testing a deep neural network-based ADAS on different pairs of three widely used simulators, demonstrates that MultiSim outperforms single-simulator testing by achieving on average a higher rate of simulator-agnostic failures by 51%. Compared to a state-of-the-art multi-simulator approach that combines the outcome of independent test generation campaigns obtained in different simulators, MultiSim identifies 54% more simulator-agnostic failing tests while showing a comparable validity rate.
To avoid the costly execution of test inputs on which simulators disagree,
we propose in addition an enhancement of MultiSim that leverages surrogate models to predict simulator disagreements and bypass executions. Our results show that utilizing a surrogate model during the search does not only increase the average number of valid failures but also improves its efficiency in finding the first valid failure. These findings indicate that combining an ensemble of simulators during the search is a promising approach for the automated cross-replication in ADAS testing.
@@ 1 Introduction
Advanced Driver Assistance Systems (ADAS) rely on perception systems like cameras and LiDAR, powered by deep neural networks (DNNs), for real-time tasks such as lane-keeping, object detection and image segmentation. While effective, these systems must operate reliably across diverse environments, yet training data cannot cover all possible scenarios survey-lei-ma . Consequently, ADAS may encounter unseen inputs post-deployment, making DNNs highly sensitive to variations in road shapes, lighting, noise, and perspective shifts. These discrepancies can lead to ADAS prediction errors, misclassifications, and inaccurate segmentations, which may impact vehicle decision-making and compromise safety Cerf:2018:CSC:3181977.3177753 .
Validating the safety of ADAS through virtual testing with scenario-based simulation is the default option for companies survey-lei-ma ; Cerf:2018:CSC:3181977.3177753 . Simulators enable developers to quickly prototype ADAS and evaluate them across a wide range of challenging scenarios. In the literature, researchers have proposed automated testing techniques to expose failing conditions and corner cases asfault ; Abdessalem-ICSE18 ; 2020-Riccio-FSE ; nejati2023reflections ; Kluck19Nsga2ADAS ; matinnejad2015 ; Zeller17SBST ; marko2019 , using various open-source ADAS simulators, such as CARLA carla , LGSVL rong2020lgsvl , BeamNG beamng , and Udacity udacity-simulator , or commercial close-source solutions, such as Siemens PreScan prescan , ESI Pro-SiVIC pro-sivic , and PTV VISSIM vissim .
However, the result of a test execution through a simulation environment only approximates the actual test outcome in the real world. Indeed, multiple studies have shown that simulation-based testing might produce inconsistent results both within-simulator, i.e., when the same test scenario is executed multiple times on the same simulator AminiFlaky2024 ; birchler2023machine , and across simulators Borg21CrossSimTesting ; matteoMaxibon2024 . This phenomenon is known as test flakiness, and it is characterized by multiple runs of the same test exhibiting non-deterministic behavior (i.e., tests pass or fail non-deterministically). Test flakiness has been studied extensively in the software testing literature parry2022survey . A study by Amini et al. AminiFlaky2024 shows that test flakiness is quite common also for simulation-based testing of ADAS, potentially leading to a distrust in virtual testing, as also reported in the field of robotic simulation icst-survey-robotics . Possible causes of flakiness stem from uncertainties in the simulation environment to timing and synchronization in the interaction between the ADAS and the simulator AminiFlaky2024 . In general, such sources of non-determinism are unknown and difficult to control, making the results of automated testing techniques, i.e., the failure-inducing test cases, unreliable.
@@
In this paper we present a novel approach named MultiSim, which incorporates multiple simulators as an ensemble directly as part of the test generation process. In particular, we cast the testing problem as a multi-objective optimization problem, where we define, for each simulator, a fitness function that evaluates the quality of the ADAS under test. By minimizing the fitness values obtained from executing each test case in multiple simulators and minimizing the distance between these values, we direct the search towards failure-inducing tests, while simultaneously generating simulator-agnostic, i.e, reliable, tests since the respective outcomes on the two simulators are close to each other.
In our evaluation of MultiSim we used the state-of-the-art DNN lane-keeping model Nvidia DAVE-2 nvidia-dave2 as system under test, and three different and widely used simulators, namely Udacity, Donkey and BeamNG asfault ; 2021-Jahangirova-ICST ; 2020-Riccio-FSE ; 2020-Stocco-ICSE ; zohdinasab2021deephyperion ; AminiFlaky2024 . In particular, we compare MultiSim with a single simulator approach, as well as with the DSS framework in terms of effectiveness, i.e., the number of simulator-agnostic failures each approach triggers given a fixed search budget, and efficiency, i.e., how quickly each approach generates the first simulator-agnostic failure.
@@
  • Approach. A novel search-based test generation approach called MultiSim that integrates multiple simulators to generate simulator-agnostic failures for ADAS.
  • Evaluation. An extensive empirical study using three simulators and three different approaches showing that MultiSim is effective at generating simulator-agnostic failures for a state-of-the-art lane-keeping ADAS. To encourage open research, our test generation approach and experimental data are available replication-package .
@@ 2.1 Simulation-based ADAS Testing
We consider Level 2 ADAS, as classified by the National Highway Traffic Safety Administration (NHTSA) nhtsa , which perform vision-based perception tasks using data gathered by camera sensors of a vehicle. Particularly, in this paper we focus on ADAS that learn the lane-keeping functionality, critical component for the safe operation of self-driving vehicles, from human-labeled driving data.
@@ 4 Approach
In the following, we describe our test case generation approach for ADAS using an ensemble of simulators. At the end of this section, we outline our validation approach to identify simulator-agnostic failures.
@@ 4.1 Ensemble-based Test Case Generation
Our proposed MultiSim approach incorporates the evaluation capabilities of multiple simulators to identify simulator-agnostic failures. The key idea is to apply ensemble learning within search-based ADAS testing, harnessing the strengths and compatibilities of individual simulators within a unified processing unit. In particular, to identify failure-inducing test inputs, MultiSim casts the testing problem as a multi-objective optimization problem, where evaluation results of multiple simulators represent objectives to be optimized. Instead of evaluating a test input in one simulator, MultiSim combines the test input evaluation results of multiple simulators in one evaluation vector. Our approach aims to obtain more reliable, and hence more valid, failure-inducing test inputs than a single-simulator approach that only evaluates a test input in a single simulator.
Our test case generation algorithm is inspired by the NSGA-II-D algorithm Sorokin2024 , which builds upon the well-established NSGA-II algorithm Deb02NSGA2 while incorporating elements of novelty search Lehman2011NoveltySearch . To enhance diversity, the algorithm maintains an archive of previously discovered test cases, optimizing the distance between new and existing tests. Additionally, it leverages a re-population operator, originally introduced in DeepJanus 2020-Riccio-FSE , which replaces a subset of dominated test inputs in the population with newly generated ones, fostering continuous exploration of the input space.
@@
1 init(T); $C\leftarrow\emptyset$; $A\leftarrow\emptyset$
@@ 4       Par $\leftarrow$ tournamentSelection$(P_{i})$
7
8      $P_{i}$leftarrow survival$(P_{i},O)$
9       $P_{i}$leftarrow repopulate$(P_{i},D,T)$
10for $\mathbf{x}\in T$ do
@@ 4.1.1 Multi-Simulator Evaluation
$A$: Archive of test inputs.
@@ 1 $Q\leftarrow\emptyset$
4
      $v_{d}\leftarrow\frac{\sum\limits_{k=1}^{j}\sum\limits_{l=k+1}^{j}|v_{k}-v_{l}|}{\binom{j}{2}}$
@@ 6       $Q\leftarrow Q\cup\{x\}$
While optimizing the functions $f_{1}(\mathbf{x}),\ldots,f_{j}(\mathbf{x})$ prioritizes failure-revealing test inputs, minimizing $f_{d}(\mathbf{x})$ drives the search towards test inputs on which all simulators agree. The fitness function $f_{a}(\mathbf{x},A)$ aims to diversify the search by calculating the distance of a test input to previously generated tests in the archive $A$ 2020-Riccio-FSE . As distance metric we the Euclidean distance between normalized test inputs. In particular, we apply Min-Max normalization, scaling values in the test input representation to a range from 0 to 1. By maximizing $f_{a}(\mathbf{x},A)$ the goal is to identify more diverse test cases. After the evaluation of a test input is completed, it is added to the archive in case its distance to the closest test in the archive exceeds a predefined threshold $\delta$ (the function addToArchive in Line 8, takes care of comparing the current test input $\mathbf{x}$ with those in the archive $A$).
@@ 4.1.2 Test Representation
In the ADAS case study, a road is defined by so-called control points, which are used to perform Catmull Spline interpolation to obtain a smooth curve 2025-Ali-ICSEW . To be able to generate diverse and challenging roads we parameterize the road definition using the vector $\mathbf{x}=(\alpha_{1},\ldots,\alpha_{n},l_{1},\ldots,l_{n})$, where $a_{i}\in[-180;180)$ denotes the clockwise angle between the i-th segment and the horizontal axis, while $l_{i}\in\mathbb{R}^{+}$ represents the length of the i-th segment between the control points $c_{i+1}$ and $c_{i+2}$. For instance, the road in 1(a) has 7 control points, 5 segments and is represented by the vector (90,110,173,120,140,10,(-90,-110,173,-120,-140,10,( - 90 , - 110 , 173 , - 120 , - 140 , 10 ,
20,18,17,15)20,18,17,15)20 , 18 , 17 , 15 ).
@@ 4.1.4 Fitness Function
For the single evaluation of a test input in one simulator we use as fitness function the function $f_{\textit{XTE}}$, which returns the maximal distance between the center of the car and the center of the lane (i.e., the cross-track error or XTE for short) over the whole execution time of the test input. We selected this function because it has been shown to be an effective fitness function for triggering failure-inducing test inputs matteoMaxibon2024 ; Humeniuk24Reinforcement ; 2020-Riccio-FSE . As oracle function we use $O=f_{\textit{XTE}}>d$ which evaluates a scenario execution as a failure when the maximal XTE value exceeds $d$.
@@ 4.2 Test Validation
An example of a feature map is shown in Figure 3. A cell with the coordinates $(x,y)$ contains test inputs whose curvature is in the range $[x-0.02,\ x+0.02]$, where $0.02$ is the granularity of the map, and the number of turns is $y$. For instance, the road in 1(a) is to be placed in the cell $(0.24,3)$ as it has a maximum of curvature 0.23 and three turns. A cell of a feature map is colored based on the road with the best fitness placed in the cells, which is ranging from green (worst, $f_{\textit{XTE}}=0$) to red (best, $f_{\textit{XTE}}=-3$) fitness value. We deem a cell as failing cell, when at least one test in the cell is failing w.r.t. the given oracle $O$. For instance, the cell $(0.32,3)$ is a failing cell, while the cell at the position $(0.04,2)$ is non-failing cell.
Figure 4: Validation of failure-inducing test inputs, i.e., $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$. Each test is re-executed in multiple simulators, i.e., two in this case $S_{1}$ and $S_{2}$, and the failure rates for each test is computed. Then the failure rates on $S_{1}$ and $S_{2}$ are compared, and the test inputs are filtered according to a failure rate threshold. (in this case 100%). In this example, the only simulator-agnostic/valid failure-inducing test input is $\mathbf{x}_{3}$.
@@ 5.1 Research Questions
The second research question allows us to evaluate whether MultiSim is efficient in identifying failures. Test efficiency is particularly relevant when the testing budget is limited Raja16NeuralNSGA2 .
@@ 5.2 Baselines
We compare the performance of MultiSim to a search-based testing approach which only employs one simulator (SingleSim) during the search.
We also compare with the recently proposed DSS approach matteoMaxibon2024 , which employs multiple simulators by re-executing test inputs generated for one simulator in a different simulator and vice-versa. DSS is, to the best of our knowledge, the only existing approach in the literature that employs multiple simulators for automated test case generation.
@@ 5.3 Procedure and Metrics
To answer all three research questions for our case study, we first validate the failures as described in Section 4.2. Specifically, we assign all tests to cells of the feature map, by extracting curvature and number of turns of the corresponding roads, using a similar feature space granularity as proposed by Biagiola et al. matteoMaxibon2024 . We randomly select three failure-inducing test inputs from each failing cell. Preliminary experiments with a higher number of selected tests per failing cell, i.e., five, did not affect the results significantly.
RQ1. To answer RQ1, to validate failures found by MultiSim, we use the third simulator that was not used during the search. We then re-execute each test five times to account for the flakiness of the simulation environment. As failure rate threshold we use 100%, i.e., a failure-inducing test input found by MultiSim is only valid if it always fails in a third simulator.
@@
[…]
RQ2. To answer RQ2, after validating all the selected failures, we identify at which time during the search each valid failure is encountered. We then select the first valid and failure-inducing test found, and compute the corresponding percentage of the search budget to find such failure.
Specifically for DSS, as it consists of two independent search executions and two migration steps, we first identify in which simulation run the failure is detected. Based on the order of the run executions and the evaluation number we compute the total run time to find the failure. That is, when a failure is detected in simulator DS1 with in total $A$ test evaluations at the evaluation number $a$, while simulator DS2 performed in total B test evaluations, we assign to the test the final evaluation count as $A+B+a$. In case the valid failure is detected during the execution of $B$ at evaluation number $b$, we assign to the test the evaluation count $A+B+A+b$, respectively.
RQ3. To answer RQ3, we select the configuration of MultiSim with the best results in terms of validity rate and number of failures, and collect test cases from past executions where the simulators both agree or disagree. In addition, we execute a modified version of MultiSim with five different seeds where the fitness function prioritizes disagreements. Specifically, we reuse Algorithm 2, but maximize $f_{d}(\mathbf{x})$, i.e., the distance between maximum XTE evaluations. We pass the roads as inputs and the labels representing an agreement or disagreement to five different and widely used ML classification models for training. Specifically, we use Decision Trees breiman1984classification , Random Forests breimannRandomF2001 , Support Vector Machines cortesSvm95 , Logistic Regression cabrera1994logistic and Stochastic Gradient Boosting FRIEDMAN2002367 . For the training of the classification models we use default hyperparameter configurations available in the Sklearn library.
We then evaluate the ML models based on F1-score, Precision, AUC-score, Accuracy. The best performing ML model on all metrics is then integrated into MultiSim and executed with the same search configuration as MultiSim.
In terms of metrics, we measure the number of valid failures (n_valid) and validity rate (valid_rate) as in RQ1, and compare MultiSim with and without the integration of a disagreement predictor. We also measure the first valid failure as the efficiency metric for both approaches.

### v2

# Simulator Ensembles for Trustworthy Autonomous Driving Systems Testing

@@ Abstract
Our empirical study, which involves testing three lane-keeping ADAS with increasing complexity on different pairs of three widely used simulators, demonstrates that MultiSim outperforms single-simulator testing by achieving, on average, a higher rate of simulator-agnostic failures by 66%. Compared to a state-of-the-art multi-simulator approach that combines the outcome of independent test generation campaigns obtained in different simulators, MultiSim identifies on average up to $3.4\times$ more simulator-agnostic failing tests and higher failure rates.
To avoid the costly execution of test inputs on which simulators disagree, we propose an enhancement of MultiSim that leverages surrogate models to predict simulator disagreements and bypass test executions. Our results show that utilizing a surrogate model during the search does not only retain the average number of valid failures but also improves its efficiency in finding the first valid failure. These findings indicate that combining an ensemble of simulators during the search is a promising approach for the automated cross-replication in ADAS testing.
@@ 1 Introduction
Advanced Driver Assistance Systems (ADAS) rely on perception systems like cameras and LiDAR, powered by deep neural networks (DNNs), for real-time tasks such as lane-keeping, object detection and image segmentation. While effective, these systems must operate reliably across diverse environments, yet training data cannot cover all possible scenarios survey-lei-ma; lou2022testing. Consequently, ADAS may encounter unseen inputs post-deployment, making DNNs highly sensitive to variations in road shapes, lighting, and noise. These discrepancies can lead to ADAS prediction errors, misclassifications, and inaccurate segmentations, which impact vehicle decision-making and compromise safety Cerf:2018:CSC:3181977.3177753; 2025-Lambertenghi-ICST.
Validating the safety of ADAS through virtual testing with scenario-based simulation is the default option for companies survey-lei-ma; Cerf:2018:CSC:3181977.3177753. Simulators enable developers to quickly prototype ADAS and evaluate them across a wide range of challenging scenarios. In the literature, researchers have proposed automated testing techniques to expose failing conditions and corner cases asfault; Abdessalem-ICSE18; 2020-Riccio-FSE; nejati2023reflections; Kluck19Nsga2ADAS; matinnejad2015; Zeller17SBST; marko2019, using various open-source ADAS simulators, such as CARLA carla, BeamNG beamng, and Udacity udacity-simulator, or commercial close-source solutions, such as Siemens PreScan prescan, ESI Pro-SiVIC pro-sivic, and PTV VISSIM vissim.
However, the result of a test execution through a simulation environment only approximates the actual test outcome in the real world. Indeed, multiple studies have shown that simulation-based testing might produce inconsistent results both within simulator, i.e., when the same test scenario is executed multiple times on the same simulator AminiFlaky2024; birchler2023machine, and across simulators Borg21CrossSimTesting; matteoMaxibon2024. This phenomenon may be due to test flakiness, and it is characterized by multiple runs of the same test exhibiting non-deterministic behavior (i.e., tests pass or fail non-deterministically). Test flakiness has been studied extensively in the software testing literature parry2022survey, including ADAS testing. A study by Amini et al. AminiFlaky2024 shows that test flakiness is quite common also for simulation-based testing of ADAS, potentially leading to a distrust in virtual testing, as also reported in the field of robotic simulation icst-survey-robotics. Possible causes of flakiness stem from uncertainties in the simulation environment to timing and synchronization in the interaction between the ADAS and the simulator AminiFlaky2024. In general, such sources of non-determinism are unknown and difficult to control, making the results of automated testing techniques, i.e., the failure-inducing test cases, unreliable.
@@
In this paper we present a novel approach named MultiSim, which incorporates multiple simulators as an ensemble directly as part of the test generation process. In particular, we cast the testing problem as a multi-objective optimization problem, where we define, for each simulator, a fitness function that evaluates the quality of the ADAS under test. In this way, we treat simulator disagreement as a first-class citizen during the optimization process. By minimizing the fitness values obtained from executing each test case in multiple simulators and their distance, we direct the search towards failure-inducing tests that are simulator-agnostic, i.e, reliable, tests since the respective outcomes on the two simulators are close to each other.
We evaluate our MultiSim approach with three lane-keeping ADAS based on convolutional DNNsnvidia-dave2, transformers 2025-Baresi-ICSE and end-to-end trajectory/control prediction tcp, on three different and widely used simulators, namely Udacity, Donkey and BeamNG asfault; 2021-Jahangirova-ICST; 2020-Riccio-FSE; 2020-Stocco-ICSE; zohdinasab2021deephyperion; AminiFlaky2024. In particular, we compare MultiSim with a single simulator approach, as well as with the DSS framework in terms of effectiveness, i.e., the number of simulator-agnostic failures each approach triggers given a fixed search budget, and efficiency, i.e., how quickly each approach generates the first simulator-agnostic failure.
@@
  • Approach. A novel search-based test generation approach called MultiSim that integrates multiple simulators to identify simulator-agnostic failures for ADAS.
  • Evaluation. An extensive empirical study using three simulators and three different approaches showing that MultiSim is effective at generating simulator-agnostic failures for three state-of-the-art lane-keeping ADAS. To encourage open research, our test generation approach and experimental data are available replication-package.
@@ 2.1 Simulation-based ADAS Testing
We consider Level 2 ADAS, as classified by the National Highway Traffic Safety Administration (NHTSA) nhtsa, which perform vision-based perception tasks using data gathered by camera sensors of a vehicle. Despite the adoption of Level 2 ADAS in many commercial vehicles, their reliability remains a concern, as evidenced by numerous recent crash reports NHTSA-level2-crashes and large scale real-world validation experiments Opletal2025ChinaADAS. Although Levels 3 and 4 ADAS have been proposed baiduapolloscapes, their real-world deployment remains highly constrained. Consequently, addressing the limitations of Level 2 systems is crucial for advancing to higher levels of autonomy. Particularly, in this paper we focus on ADAS that learn the lane-keeping functionality, critical component for the safe operation of self-driving vehicles, from human-labeled driving data.
@@ 4.1 Ensemble-based Test Case Generation
Our proposed MultiSim approach incorporates the evaluation capabilities of multiple simulators to identify simulator-agnostic failures. The key idea is to apply ensemble learning within search-based ADAS testing, harnessing the strengths and compatibilities of individual simulators within a unified processing unit. We reformulated the multi-objective optimization problem to incorporate simulator-specific agreement and disagreement objectives, allowing the search to explicitly reason about the consistency of failures across heterogeneous simulators. Instead of evaluating a test input in one simulator, MultiSim combines the test input evaluation results of multiple simulators in one evaluation vector. Our approach aims to obtain more reliable, and hence more valid, failure-inducing test inputs than a single-simulator approach that only evaluates a test input in a single simulator.
@@
1
21exinit(T); $C\leftarrow\emptyset$; $A\leftarrow\emptyset$
@@ 3 $P_{0}\leftarrow\textsc{randomSampling}(T,D)$
8
[…]
11for $\mathbf{x}\in T$ do
@@ 4.1.1 Multi-Simulator Evaluation
AA: Archive of test inputs.
@@ 1 $Q\leftarrow\emptyset$
5
 $v_{d}\leftarrow\frac{\sum\limits_{k=1}^{j-1}\sum\limits_{l=k+1}^{j}|v_{k}-v_{l}|}{\binom{j}{2}}$
@@ 7  $Q\leftarrow Q\cup\{x\}$
While optimizing the functions $f_{1}(\mathbf{x}),\ldots,f_{j}(\mathbf{x})$ prioritizes failure-revealing test inputs, minimizing $f_{d}(\mathbf{x})$ drives the search towards test inputs on which all simulators agree. The fitness function $f_{a}(\mathbf{x},A)$ aims to diversify the search by calculating the distance of a test input to previously generated tests in the archive $A$ 2020-Riccio-FSE. As distance metric we use the Euclidean distance between normalized test inputs. In particular, we apply Min-Max normalization, scaling values in the test input representation to a range from 0 to 1. By maximizing $f_{a}(\mathbf{x},A)$ the goal is to identify more diverse test cases. After the evaluation of a test input is completed, it is added to the archive in case its distance to the closest test in the archive exceeds a predefined threshold $\delta$ (the function addToArchive in Line 8, takes care of comparing the current test input $\mathbf{x}$ with those in the archive $A$).
@@ 4.1.2 Test Representation
We adopt a model-based representation of test inputs. In line with previous studies 2020-Riccio-FSE; 2025-Ali-ICSEW, each road scenario is defined through a sequence of control points. These points are interpolated using Catmull–Rom splines to generate smooth and continuous trajectories. To be able to generate diverse and challenging roads we parameterize the road definition using the vector $\mathbf{x}=(\alpha_{1},\ldots,\alpha_{n},l_{1},\ldots,l_{n})$, where $a_{i}\in[-180;180)$ denotes the clockwise angle between the i-th segment and the horizontal axis, while $l_{i}\in\mathbb{R}^{+}$ represents the length of the i-th segment between the control points $c_{i+1}$ and $c_{i+2}$. For instance, the road in 2(a) has 7 control points, 5 segments and is represented by the vector (90,110,173,120,140,10,20,18,17,15)(-90,-110,173,-120,-140,10,20,18,17,15). Note that this choice of input representation is not exclusive of our approach and equivalent parameterized representations can be defined, following similar principles 2025-Chen-arxiv.
@@ 4.1.4 Fitness Function
For the evaluation of a single test input within one simulator, we define a fitness function that captures the system-level behavior of the SUT. In the case of lane-keeping ADAS, the chosen fitness function is the cross-track error, denoted as $f_{\textit{XTE}}$. This function computes the maximum lateral deviation between the vehicle’s center and the lane center over the entire duration of the test execution, providing a direct quantitative measure of lane-keeping performance. We selected this function because it has been shown to be an effective fitness function for triggering failure-inducing test inputs matteoMaxibon2024; Humeniuk24Reinforcement; 2020-Riccio-FSE. As oracle function we use $O=f_{\textit{XTE}}>d$ which evaluates a scenario execution as a failure when the maximal XTE value exceeds $d$.
@@ 4.2 Test Validation
An example of a feature map is shown in Figure 3. A cell with the coordinates $(x,y)$ contains test inputs whose curvature is in the range [x0.02,x+0.02][x-0.02,\ x+0.02], where $0.02$ is the granularity of the map, and the number of turns is $y$. For instance, the road in 1(a) is to be placed in the cell $(0.24,3)$ as it has a maximum of curvature 0.23 and three turns. A cell of a feature map is colored based on the road with the best fitness placed in the cells, which is ranging from green (worst, $f_{\textit{XTE}}=0$) to red (best, $f_{\textit{XTE}}=-3$) fitness value. We deem a cell as failing cell, when at least one test in the cell is failing w.r.t. the given oracle $O$. For instance, the cell $(0.32,3)$ is a failing cell, while the cell at the position $(0.04,2)$ is non-failing cell.
Figure 4: Validation of failure-inducing test inputs, i.e., $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$. Each test is re-executed in multiple simulators, i.e., two in this case $S_{1}$ and $S_{2}$, and the failure rates for each test are computed. Then the failure rates on $S_{1}$ and $S_{2}$ are compared, and the test inputs are filtered according to a failure rate threshold. (in this case 100%). In this example, the only simulator-agnostic/valid failure-inducing test input is $\mathbf{x}_{3}$.
@@ 5.1 Research Questions
The second research question evaluates whether MultiSim is efficient in identifying failures. Test efficiency is particularly relevant when the testing budget is limited Raja16NeuralNSGA2.
@@ 5.2 Baselines
We compare the performance of MultiSim to a search-based testing approach which only employs one simulator (SingleSim) during the search. We also compare with the recently proposed DSS approach matteoMaxibon2024, which employs multiple simulators by re-executing test inputs generated for one simulator in a different simulator and vice-versa. DSS is, to the best of our knowledge, the only existing approach in the literature that employs multiple simulators for automated test case generation in the ADAS domain.
@@ 5.3 Procedure and Metrics
To answer all three research questions for our experiment, we first validate the failures as described in Section 4.2. Specifically, we assign all tests to cells of the feature map, by extracting curvature and number of turns of the corresponding roads, using the same feature space granularity as proposed by Biagiola et al. matteoMaxibon2024. We randomly select three failure-inducing test inputs from each failing cell. Preliminary experiments with a higher number of selected tests per failing cell, i.e., five, did not affect the results significantly.
To answer RQ1, to validate failures found by MultiSim, we use the third simulator that was not used during the search. We then re-execute each test five times to account for the flakiness of the simulation environment. As failure rate threshold we use 100%, i.e., a failure-inducing test input found by MultiSim is only valid if it always fails in the third validation simulator.
@@
$\text{valid\_rate}=\frac{\text{\# valid failures}}{\text{\# failures selected for validation}}$
To answer RQ2, after validating all the selected failures, we identify at which time during the search each valid failure is encountered. Thus, we define efficiency as the proportion of the search budget required to identify the first valid failure. It measures how early in the search process the approach discovers a fault-revealing test relative to the total number of generations or evaluations. This metric reflects the practical usefulness of the search procedure in time- and resource-constrained testing contexts, where faster identification of valid failures is beneficial.
For DSS, as it consists of two independent search executions and two migration steps, we first identify in which simulation run the failure is detected. Based on the order of the run executions and the evaluation number we compute the total run time to find the failure. That is, when a failure is detected in simulator DS1 with in total $A$ test evaluations at the evaluation number $a$, while simulator DS2 performed in total B test evaluations, we assign to the test the final evaluation count as $A+B+a$. In case the valid failure is detected during the execution of $B$ at evaluation number $b$, we assign to the test the evaluation count $A+B+A+b$, respectively. We evaluate efficiency using this methodology because DSS consists of two independent testing experiments, and unlike single-simulator-based testing, it is not possible to incrementally order the test cases generated by DSS for a single run.
To answer RQ3, we select the configuration of MultiSim with the best results in terms of validity rate and number of failures, and collect test cases from past executions where the simulators both agree or disagree. In addition, we execute a modified version of MultiSim with five different seeds where the fitness function prioritizes disagreements. Specifically, we reuse Algorithm 2, but maximize $f_{d}(\mathbf{x})$, i.e., the distance between maximum XTE evaluations. We pass the roads as inputs and the labels representing an agreement or disagreement to five different and widely used ML classification models for training. Specifically, we use Decision Trees breiman1984classification, Random Forests breimannRandomF2001, Support Vector Machines cortesSvm95, Logistic Regression cabrera1994logistic and Stochastic Gradient Boosting FRIEDMAN2002367. For the training of the classification models we use default hyperparameter configurations available in the Sklearn library.
We then evaluate the ML models based on F-1-score and AUC-score. The best performing ML model on all metrics is then integrated into MultiSim and executed with the same search configuration as MultiSim.
In terms of metrics, we measure the number of valid failures (n_valid) and validity rate (valid_rate) as in RQ1, and compare MultiSim with and without the integration of a disagreement predictor. We also measure the first valid failure as the efficiency metric for both configurations.